# XATU: A Fine-grained Instruction-based Benchmark for Explainable Text Updates

### v1

@@ Abstract
Text editing is a crucial task that involves modifying text to better align with user intents. However, existing text editing benchmark datasets have limitations in providing only coarse-grained instructions. Consequently, although the edited output may seem reasonable, it often deviates from the intended changes outlined in the gold reference, resulting in low evaluation scores. To comprehensively investigate the text editing capabilities of large language models, this paper introduces XATU, the first benchmark specifically designed for fine-grained instruction-based explainable text editing. XATU covers a wide range of topics and text types, incorporating lexical, syntactic, semantic, and knowledge-intensive edits. To enhance interpretability, we leverage high-quality data sources and human annotation, resulting in a benchmark that includes fine-grained instructions and gold-standard edit explanations. By evaluating existing open and closed large language models against our benchmark, we demonstrate the effectiveness of instruction tuning and the impact of underlying architecture across various editing tasks. Furthermore, extensive experimentation reveals the significant role of explanations in fine-tuning language models for text editing tasks. The benchmark will be open-sourced to support reproduction and facilitate future research. [^1] The code will be available at https://github.com/megagonlabs/xatu
1 Introduction
@@
Recent advances in large language models (LLMs) have demonstrated remarkable zero-shot text generation capabilities across a wide range of downstream natural language processing (NLP) tasks like question answering, dialogue, and summarization Radford et al. (2019); Raffel et al. (2020); Brown et al. (2020); Zhang et al. (2022, 2023a). By incorporating instruction tuning Wei et al. (2022) and reinforcement learning from human feedback (RLHF) Bai et al. (2022); Ouyang et al. (2022), state-of-the-art LLMs exhibit a remarkable ability to effectively align with user instructions.
However, most such work adopts a one-shot paradigm, generating text in a single step. In contrast, human text composition often involves an iterative process consisting of multiple drafts and revisions Faltings et al. (2021). This iterative process allows for refining and improving the text over time, ensuring better alignment with the user’s intents, while the current setting of generating text contents in a single pass can be highly restrictive Hayes and Flower (1980); Collins and Gentner (1980). An alternative approach that offers more flexibility is iterative text generation and refinement, involving successive iterations of modular updates and modifications to the text Vaughan and McDonald (1986). These iterations focus on text editing tasks such as paraphrasing and updating information to enhance the overall quality of the generated output.
@@
Consequently, there has been a growing interest among researchers in instruction-based text editing tasks in recent years. Text editing plays a crucial role in enabling interactivity, as it allows users to modify previously generated text to suit their dynamic constraints. This represents a step further beyond one-shot generation towards interactive text generation. Specifically, Faltings et al. (2021) introduced the concept of controllable text editing, wherein a model learns to edit the text in response to user commands while incorporating grounding to prevent hallucination issues. Moreover, Dwivedi-Yu et al. (2022) proposed EditEval, a benchmark and evaluation suite that utilizes high-quality existing and new datasets to automatically assess the editing capabilities of different systems.
Nevertheless, these works only consider simple commands or instructions, such as ‘Update the article’ or ‘Expand’, resulting in three main limitations. (a) These instructions are often too coarse-grained, to the extent that even humans may struggle to understand the desired intent of the edit, limiting models’ instruction following capabilities. (b) It poses challenges for text editing evaluation since outputs obtained using coarse-grained instructions may not accurately reflect the true editing capabilities of LLM systems. The coarse-grained instructions cannot provide sufficient guidance for the editing system, so the edited output may appear correct but deviate from the intended changes, as shown in Figure 1. The edited output updates the death information of Isidore Mankofsky correctly but since the ground truth output updates the information about his famous works. This edit will receive very low evaluation scores since current automatic evaluation methods for text editing systems are mostly based on surface-level overlap with human-written gold references. (c) The current text editing systems suffer from a lack of explanation ability, which in turn limits the interpretability of the edits made. This lack of interpretability hampers our understanding of the underlying text editing behaviors.
To address the aforementioned limitations, we present XATU, a novel fine-grained instruction-based benchmark designed for explainable text updates. The benchmark leverages high-quality existing data sources to enable automatic evaluation of LLM editing capabilities. In comparison to other datasets and benchmarks, XATU highlights its inclusion of a wider range of diverse topics, fine-grained edit instructions, and corresponding explanation rationales, as shown in Table 1. By utilizing XATU, we are able to evaluate and compare the performance of several state-of-the-art language models, considering both zero-shot and fine-tuning settings. Experiment results emphasize the significant role of explanations in the fine-tuning process of language models for text editing tasks. Our findings shed light on the importance of incorporating explanation rationales to enhance the performance of language models in text editing, ultimately leading to improved output quality. We summarize the contributions of this paper as follows:
@@
2 Related Work
2.1 Text Editing
Several prior studies have concentrated on the domain of text editing. Guu et al. (2018) introduced a method that involves generating sentences through the editing of training prototypes. Faruqui et al. (2018) presented the WikiAtomicEdits dataset, which comprises edits extracted from Wikipedia editing history. Subsequently, Yin et al. (2019); Reid and Neubig (2022) leveraged Wikipedia editing history to propose the task of learning to represent edits. Furthermore, Iso et al. (2020) introduced a fact-based text editing task in their work.
@@ 2.1 Text Editing
2.2 Text Editing Datasets
Previous research has also introduced various datasets and resources that focus on iterative text revisions within specific domains. For instance, datasets presented in  Yang et al. (2017); Faruqui et al. (2018); Anthonio et al. (2020) primarily concentrated on the domain of Wikipedia edit history, while Spangher et al. (2022) developed a dataset tailored to news articles. Researchers in the field have also proposed more comprehensive benchmarks that cover multiple domains. IteraTeR Du et al. (2022) provides iterative tasks from multiple domains, but has only a limited number of tasks, such as fluency, coherence, clarity, style, and meaning change. EditEval Dwivedi-Yu et al. (2022) is perhaps the closest to our benchmark XATU in that it covers data from multiple domains: Wikipedia, Wikinews, news articles, and arXiv.
@@ 2.2 Text Editing Datasets
3 The XATU Benchmark
3.1 Data Source
@@
In line with the systematic formulation of natural language styles in natural languages presented by DiMarco and Hirst (1993), we adopt a similar approach to categorize text editing into four main editing aspects: lexical, syntactic, semantic, and knowledge. Our benchmark XATU aims to serve as a relatively comprehensive benchmark and a holistic evaluation framework for text editing tasks, covering all four aspects of text editing.
@@
To ensure the comprehensiveness and quality of our benchmark, we meticulously curate high-quality editing data from $4$ distinct downstream NLP tasks, encompassing a total of $1000$ annotated data from $9$ data sources, inspired by the criteria outlined in DiMarco and Hirst (1993). A summary of the tasks and datasets, along with an overview of the editing aspects covered by each dataset, is presented in Table 2. We also transform all datasets within the XATU benchmark into a consistent format, extracting the input text, gold edits, task type, and optional reference documents from the following data sources, as illustrated in the example in Figure 2. This standardized format facilitates evaluation and comparison across different datasets within the benchmark.
{tcolorbox}
@@
[fontupper=, colback=green!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Instruction: Modify the input text to incorporate details about the nature of the matches played, the players included in the tour, mention of Nat Sciver captaining the England team for the first time, and the simultaneous scheduling of New Zealand men’s matches. {tcolorbox}[fontupper=, colback=blue!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Input: The England women’s cricket team is scheduled to play against New Zealand women’s cricket team in February and March 2021. The tour is scheduled to be played during the time that was originally scheduled to be used to host the … {tcolorbox}[fontupper=, colback=blue!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Reference: (0) The last three T201 matches took place on the same day as the New Zealand women’s fixtures at the same venues…
@@
(7) Sophie_Ecclestone International career In February 2021, during England’s tour of New Zealand, Ecclestone took her 100th international wicket.
(8) Sarah_Glenn International career In 2021, Glenn was named in the squad for England’s tour of New Zealand.
@@
{tcolorbox}[fontupper=, colback=blue!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Output: The England women’s cricket team played against the New Zealand women’s cricket team in February and March 2021. The series was played during the time that was originally scheduled to be used to host the 2021 Women’s Cricket World Cup, after the tournament was postponed by one year due to the COVID-19 pandemic… {tcolorbox}[fontupper=, colback=green!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Explanation: The text editing system was instructed to update the text based on the references provided. It corrected the phrasing to past tense, indicating that the England and New Zealand women’s cricket teams already ’played’ the matches instead of ’scheduled to play’. The information about the tour, including the confirmation of fixtures…
@@
We first include the Wiki Neutrality Corpus (WNC) Pryzant et al. (2020), a collection of original and de-biased sentence pairs mined from Wikipedia edits by carefully filtering based on the editor’s comments.
We also adopt data from Wikibias Zhong (2021), a manually annotated parallel corpus with sentence pairs from Wikipedia editing history. The inclusion of Wikibias complements the mined data from WNC, providing additional annotations to address both sentence-level and token-level biases.
Furthermore, we also include data from StylePTB Lyu et al. (2021). StylePTB contains paired sentences undergoing various fine-grained stylistic changes and compositions of multiple transfers.
@@
We include the FRUIT dataset (Iv et al., 2022), a dataset collected by comparing two snapshots of the same Wikipedia article. The reference documents were identified by searching for other Wikipedia articles and human filtering. We include this gold set from the FRUIT dataset in XATU.
@@
DeFacto Liu et al. (2022) dataset contains consists of document summaries from the XSum dataset Narayan et al. (2018), accompanied by human-corrected versions that rectify factual errors present in the original summaries. For the text editing task, we utilize the original document as the reference to evaluate the ability of text editing systems to correct factual errors in the summaries effectively. The inclusion of this dataset allows us to assess the performance of systems in editing text to ensure factual accuracy, an important aspect of text editing in real-world applications.
@@
3.2 Annotation Process
Serving as a fine-grained instruction-based and explainable benchmark, each instance in XATU comprises five key components as shown in Figure 2: inputs, outputs, optional references, fine-grained instructions, and explanations. Since the source datasets lacked fine-grained instructions and explanations, we meticulously annotated them.
@@ 3.2 Annotation Process
Annotating fine-grained instructions and explanations poses a challenge due to the inherently open-ended nature of the text production task, making it difficult to filter out low-quality workers. Additionally, many workers turn to LLMs for text production tasks due to their prevalence Veselovsky et al. (2023), but the output quality heavily depends on the user’s LLM skills Zamfirescu-Pereira et al. (2023). Therefore, relying solely on crowd workers for annotation may not be advisable.
To address these concerns and maintain the quality and reliability of the annotated data, we implemented an LLM-in-the-loop annotation approach, which involves generating candidates of fine-grained instruction and explanations using LLMs and then validating the output with human workers. This allowed us to mitigate the challenges associated with crowd-sourcing and refine our methodology for generating and evaluating fine-grained instructions and edit explanations.
{tcolorbox}
[fontupper=, fontlower=, colback=gray!5, left=1mm, right=1mm, top=1mm, bottom=1mm] Input: At least once an episode we see protestors marching around screaming slogans
[…]
Output: <ins>One or more times</ins> an episode protestors <ins>are seen walking</ins> around <ins>yelling blue</ins> slogans <ins>.</ins>
@@
We generate candidates of fine-grained instructions and explanations for each instance using LLMs Zhou et al. (2023b); Honovich et al. (2023). For the LLMs to generate candidates, we use GPT-4 OpenAI (2023). Through our preliminary prompt engineering, we found that merely asking LLMs to generate the instructions and explanations often resulted in undesired outputs [^2] More details can be found in the Appendix.. We found that adding HTML tags to explicitly indicate the edited portion is highly effective, as illustrated in Figure 3. Thus, we decided to use this prompting technique to generate instructions and explanations. The full prompt can be found in the Appendix.
@@ 3.3 Benchmark Usage
4 Experiments
In this section, we perform a series of text editing experiments using the XATU benchmark. These experiments aim to investigate and address the following research questions:
@@ 4 Experiments
By evaluating the performance of LLMs, we aim to assess their effectiveness in carrying out text editing tasks and understand their limitations and strengths in this context.
  • Q2: How do our fine-grained instructions differ from the coarse-grained instructions that are commonly provided in text editing tasks?
@@
4.1 Experimental Setup
We include the following baseline large languages models in our experiments:
@@ 4.1 Experimental Setup
We use the Huggingface implementation of all models with LoRA Hu et al. (2022) for the fine-tuning experiments. We instruction-tuned all models for $15$ epochs with $200$ training examples at a learning rate of $51e-$ 5 and tested on the original $1000$ test set, as shown in Table 2. We keep a small training data size as Zhou et al. (2023a) found that alignment doesn’t require too much data. The detailed training parameters are presented in Section B.
4.2 Zero-shot Results

### v2

@@ Abstract
Text editing is a crucial task of modifying text to better align with user intents. However, existing text editing benchmark datasets contain only coarse-grained instructions and lack explainability, thus resulting in outputs that deviate from the intended changes outlined in the gold reference. To comprehensively investigate the text editing capabilities of large language models (LLMs), this paper introduces XATU, the first benchmark specifically designed for fine-grained instruction-based explainable text editing. XATU considers finer-grained text editing tasks of varying difficulty (simplification, grammar check, fact-check, etc.), incorporating lexical, syntactic, semantic, and knowledge-intensive edit aspects. To enhance interpretability, we combine LLM-based annotation and human annotation, resulting in a benchmark that includes fine-grained instructions and gold-standard edit explanations. By evaluating existing LLMs against our benchmark, we demonstrate the effectiveness of instruction tuning and the impact of underlying architecture across various editing tasks. Furthermore, extensive experimentation reveals the significant role of explanations in fine-tuning language models for text editing tasks. The benchmark will be open-sourced to support reproduction and facilitate future research at https://github.com/megagonlabs/xatu.
Keywords: Text Editing, Large Language Model, Language Resources, Explainability
[…]
XATU: A Fine-grained Instruction-based
Benchmark for Explainable Text Updates
[…]
haopeng@ifmlab.org
{hayate, sairam, nikita}@megagon.ai
Abstract content
The work was done when Haopeng Zhang was a research intern at Megagon Labs
1.   Introduction
Text editing is the task of modifying text to better align with user intents. Recent advances in large language models (LLMs) have demonstrated remarkable zero-shot text generation capabilities across a wide range of downstream natural language processing (NLP) tasks like question answering, dialogue, and summarization Radford et al. (2019); Raffel et al. (2020); Brown et al. (2020); Zhang et al. (2022, 2023a). It has been shown that incorporating instruction tuning Wei et al. (2022) and reinforcement learning from human feedback (RLHF) Bai et al. (2022); Ouyang et al. (2022) can further enhance a model’s ability to align with the user’s intent.
Text editing plays a crucial role in real-world text generation applications. While many current text generation models typically employ a one-shot manner, humans often engage in an iterative process when writing text, which entails multiple drafts and revisions Faltings et al. (2021). Thus, approaching text generation as an iterative process with successive updates to the text is a more effective way to achieve higher alignment between the model’s outputs and the user’s intent.
@@ 1.   Introduction
Consequently, instruction-based text editing tasks have received growing interest recently. They enable a user to interact with a model with commands to update existing text and achieve desirable text. Typically commands target broad edits such as adding or removing content, or changing the meaning of text Faltings et al. (2021). To evaluate models’ capabilities for text editing, instruction-based benchmarks such as EditEval Dwivedi-Yu et al. (2022) have been proposed recently.
Nevertheless, prior datasets and benchmarks only consider simple commands or instructions, such as ‘Update the article’ or ‘Expand’, resulting in three main limitations. (a) These instructions are often too coarse-grained, to the extent that even humans may struggle to understand the desired intent of the edit, limiting a model’s capability to follow instructions. (b) It poses challenges for text editing evaluation since outputs obtained using coarse-grained instructions may not accurately reflect the true editing capabilities of LLM systems. The coarse-grained instructions cannot provide sufficient guidance for the editing system, so the edited output may appear correct but deviate from the intended changes, as shown in Figure 1. The edited output updates the death information of “Isidore Mankofsky” correctly but since the ground truth output updates the information about his famous works. This edit will receive very low evaluation scores since current automatic evaluation methods for text editing systems are mostly based on surface-level overlap with human-written gold references. (c) The current text editing systems lack explainability, which in turn limits the interpretability of the edits made. This lack of interpretability hampers our understanding of the underlying text editing behaviors.
@@
To address the aforementioned limitations, we present XATU, a novel fine-grained instruction-based benchmark designed for explainable text updates. The benchmark leverages high-quality existing data sources from different tasks to enable automatic evaluation of LLM editing capabilities by incorporating an LLM-in-the-loop annotation process. In comparison to other datasets and benchmarks, XATU highlights its inclusion of a wider range of diverse topics, fine-grained edit instructions, and corresponding explanation rationales, as shown in Table 1. By utilizing XATU, we are able to evaluate and compare the performance of several state-of-the-art language models, considering both zero-shot and fine-tuning settings. Experiment results emphasize the significant role of explanations in the fine-tuning process of language models for text editing tasks. Our findings shed light on the importance of incorporating explanation rationales to enhance the performance of language models in text editing, ultimately leading to improved output quality. We summarize the contributions of this paper as follows:
@@
2.   Related Work
2.1.   Text Editing
Several prior studies have concentrated on the domain of text editing. Guu et al. (2018) introduced a method that involves generating sentences through the editing of training prototypes. Faruqui et al. (2018) presented the WikiAtomicEdits dataset, which comprises edits extracted from Wikipedia editing history. Subsequently, Yin et al. (2019) and Reid and Neubig (2022) leveraged Wikipedia editing history to propose the task of learning to represent edits. Furthermore, Iso et al. (2020) introduced a fact-based text editing task in their work.
@@ 2.1.   Text Editing
2.2.   Text Editing Datasets
Previous research has also introduced various datasets and resources that focus on iterative text revisions within specific domains. For instance, datasets presented in  Yang et al. (2017),  Faruqui et al. (2018) and  Anthonio et al. (2020) primarily concentrated on the domain of Wikipedia edit history, while Spangher et al. (2022) developed a dataset tailored to news articles. Researchers in the field have also proposed more comprehensive benchmarks that cover multiple domains. IteraTeR Du et al. (2022) provides iterative tasks from multiple domains, but has only a limited number of tasks, such as fluency, coherence, clarity, style, and meaning change. EditEval Dwivedi-Yu et al. (2022) is perhaps the closest to our benchmark XATU in that it covers data from multiple domains: Wikipedia, Wikinews, news articles, and arXiv.
@@ 2.2.   Text Editing Datasets
3.   The XATU Benchmark
We introduced the XATU benchmark in this section. We outline the criteria for selecting datasets to construct XATU (§3.1), how to annotate fine-grained instructions and explanations for the edits (§3.2), and discuss various use cases (§3.3). As shown in Figure 2, each instance in XATU comprises five key components: inputs, outputs, optional references, fine-grained instructions, and explanations.
3.1.   Data Source
@@
In line with the systematic formulation of language style DiMarco and Hirst (1993), we adopt a similar approach to categorize text editing into four main editing aspects: lexical, syntactic, semantic, and knowledge. Our benchmark XATU aims to serve as a relatively comprehensive benchmark and a holistic evaluation framework for text editing tasks, covering all four aspects of text editing.
@@
To ensure comprehensiveness and quality, we meticulously curate high-quality editing data from $4$ distinct downstream NLP tasks, encompassing a total of $1000$ annotated data from $9$ data sources following DiMarco and Hirst (1993). A summary of the tasks and datasets, along with an overview of the editing aspects covered by each dataset, is presented in Table 2. We also transform all datasets within the XATU benchmark into a consistent format, extracting the input text, gold edits, task type, and optional reference documents from the following data sources, as illustrated in the example in Figure 2. This standardized format facilitates evaluation and comparison across different datasets within the benchmark.
@@
Instruction: Modify the input text to incorporate details about the nature of the matches played, the players included in the tour, mention of Nat Sciver captaining the England team for the first time, and the simultaneous scheduling of New Zealand men’s matches.
Input: The England women’s cricket team is scheduled to play against New Zealand women’s cricket team in February and March 2021. The tour is scheduled to be played during the time that was originally scheduled to be used to host the …
@@
Reference: (0) The last three T201 matches took place on the same day as the New Zealand women’s fixtures at the same venues… (7) Sophie_Ecclestone International career In February 2021, during England’s tour of New Zealand, Ecclestone took her 100th international wicket. (8) Sarah_Glenn International career In 2021, Glenn was named in the squad for England’s tour of New Zealand.
@@
Output: The England women’s cricket team played against the New Zealand women’s cricket team in February and March 2021. The series was played during the time that was originally scheduled to be used to host the 2021 Women’s Cricket World Cup, after the tournament was postponed by one year due to the COVID-19 pandemic…
Explanation: The text editing system was instructed to update the text based on the references provided. It corrected the phrasing to past tense, indicating that the England and New Zealand women’s cricket teams already ’played’ the matches instead of ’scheduled to play’. The information about the tour, including the confirmation of fixtures…
@@
  • Wiki Neutrality Corpus (WNC) Pryzant et al. (2020) is a collection of original and de-biased sentence pairs mined from Wikipedia edits by carefully filtering based on the editor’s comments.
  • Wikibias Zhong (2021) is a manually annotated parallel corpus with sentence pairs from Wikipedia editing history. The inclusion of Wikibias complements the mined data from WNC, providing additional annotations to address both sentence-level and token-level biases.
  • StylePTB Lyu et al. (2021) contains paired sentences undergoing various fine-grained stylistic changes and compositions of multiple transfers.
@@
  • FRUIT dataset (Iv et al., 2022) is a dataset collected by comparing two snapshots of the same Wikipedia article. The reference documents were identified by searching for other Wikipedia articles and human filtering. We include this gold set from the FRUIT dataset in XATU.
@@
  • DeFacto Liu et al. (2022) dataset consists of document summaries from the XSum dataset Narayan et al. (2018), accompanied by human-corrected versions that rectify factual errors present in the original summaries. For the text editing task, we utilize the original document as the reference to evaluate the ability of text editing systems to correct factual errors in the summaries effectively. The inclusion of this dataset allows us to assess the performance of systems in editing text to ensure factual accuracy, an important aspect of text editing in real-world applications.
@@
3.2.   Annotation Process
@@
Annotating fine-grained instructions and explanations poses a challenge due to the inherent open-ended nature of the text production task, making it difficult to filter out low-quality workers. Additionally, many workers turn to LLMs for text production tasks due to their prevalence Veselovsky et al. (2023), but the output quality heavily depends on the user’s LLM skills Zamfirescu-Pereira et al. (2023). Therefore, relying solely on crowd workers for annotation may not be advisable.
To address these concerns and maintain the quality and reliability of the annotated data, we implemented an LLM-in-the-loop annotation approach, which involves generating candidates of fine-grained instruction and explanations using LLMs and then validating the output with human workers. This allowed us to mitigate the challenges associated with crowd-sourcing and refine our methodology for generating and evaluating fine-grained instructions and editing explanations.
Input: At least once an episode we see protestors marching around screaming slogans Output: One or more times an episode protestors are seen walking around yelling blue slogans. Input: <del>At least once</del> an episode <del>we see</del> protestors <del>marching</del> around <del>screaming</del> slogans Output: <ins>One or more times</ins> an episode protestors <ins>are seen walking</ins> around <ins>yelling blue</ins> slogans <ins>.</ins>
@@
We generate candidates of fine-grained instructions and explanations for each instance using LLMs Zhou et al. (2023b); Honovich et al. (2023); Li et al. (2023). For the LLMs to generate candidates, we use GPT-4 OpenAI (2023). Through our preliminary prompt engineering, we found that merely asking LLMs to generate the instructions and explanations often resulted in undesired outputs [^1] More details can be found in the Appendix.. We found that adding HTML tags to explicitly indicate the edited portion is highly effective, as illustrated in Figure 3. Thus, we decided to use this prompting technique to generate instructions and explanations. The full prompt can be found in the Appendix.
@@ 3.3.   Benchmark Usage
4.   Experiments
In this section, we perform a series of text editing experiments using the XATU benchmark, aiming to investigate and address the following research questions:
@@ 4.   Experiments
By evaluating the performance of LLMs, we aim to assess their effectiveness of instruction-based text editing and understand their limitations and strengths in this context.
  • Q2: How do fine-grained instructions in XATU differ from the coarse-grained instructions?
@@
4.1.   Experimental Setup
We include the following baseline LLMs in our experiments:
@@ 4.1.   Experimental Setup
We use the Huggingface implementation of all models with LoRA Hu et al. (2022) for the fine-tuning experiments. We instruction-tuned all models for $15$ epochs with $200$ training examples at a learning rate of $51e-$ 5 and tested on the original $1000$ test set, as shown in Table 2. We keep a small training data size as Zhou et al. (2023a) found that alignment doesn’t require too much data. We use LoRA Hu et al. (2022) to instruction fine-tune all models for $15$ epochs with $200$ additional training examples at a learning rate of $51e-$ 5.
4.2.   Zero-shot Results